What is the pressure drop due to the Bernoulli Effect as water goes into a 4.44-cm-diameter nozzle from a 15.25-cm-diameter fire hose while carrying a flow of 72 L/s? (b) To what maximum height above the nozzle can this water rise? 


Round your answer to 0 decimal places


1
Expert's answer
2021-05-04T06:36:20-0400

To be given in question

Radius

R1=7.625"\times10^{-2}m"

R2=2.22"\times10^{-2}m"

"Q=72\times10^{-3}m/sec"

To be asked in question

Velocity

v1=?

v2=?

Part(a)

We know that

Bernoulli eqution

"P_{1}+\frac{1}{2}\rho v_{1}^2+\rho gh_{1}=P_{2}+\frac{1}{2}\rho v_{2}^2+\rho gh_{2}"


We know that h=h1=h2

"P_{1}+\frac{1}{2}\rho v_{1}^2=P_{2}+\frac{1}{2}\rho v_{2}^2\rightarrow(1)"

Continuity equation

"Q=A_{1}v_{1}=A_{2}v_{2}\rightarrow(2)"

"v_{1}=\frac{Q}{A_{1}}"

Put value

"v_{1}=\frac{72\times10^{-3}}{3.14\times(7.625\times10^{-2})^2}"

"v_{1}=3" m/sec

"v_{2}=\frac{Q}{A_{2}}"

Put value

"v_{2}=\frac{72\times10^{-3}}{3.14\times(2.22\times10^{-2})^2}"

"v_{2}=46m\/sec"


Part (b)

We can re -written as equation (2)

"P_{1}+\frac{1}{2}\rho v_{1}^2=P_{2}+\frac{1}{2}\rho v_{2}^2"

"∆P=\frac{1}{2}\rho v_{2}^2-\frac{1}{2}\rho v_{1}^2"

"∆P=\frac{1}{2}\rho( v_{2}^2-v_{1}^2)\rightarrow(3)"

Put value in eqution (3)

"∆P=1.053\times10^6N/m^2"